OpenStudy (anonymous):

A deck of cards consists of 8 blue cards and 5 white cards. A simple random sample (random draws without replacement) of 6 cards is selected. What is the chance that one of the colors appears twice as many times as the other?

OpenStudy (anonymous):

@agent0smith m unable to give u reply...dnt no y..??

OpenStudy (agent0smith):

We need to get either: 4 blues and 2 whites: from 8 blues, choose 4, from 5 whites, choose 2. Total is 13, so from 13 choose 6. \[\Large \frac{ 8C4 \times 5C2 }{13C6 }\] or 4 whites and 2 blues..see if you can get this one :) Then you add the two together.

OpenStudy (anonymous):

@agent0smith some server error i think..:(...

OpenStudy (agent0smith):

Yeah, i got that too :/

OpenStudy (anonymous):

i got tht 1 wrong...m gettin 0.0349...

OpenStudy (agent0smith):

Show your work for the second part, 4 whites 2 blue

OpenStudy (anonymous):

..the answer to this is 0.4895

OpenStudy (anonymous):

you need 4b 2w or 4w 2b = p 4b 2w + p 4w 2b p 4b 2w = 40.79254% p 4w 2b = 8.15851%

OpenStudy (agent0smith):

This is actually the work that those guys didn't show...\\[\Large \frac{ 8C4 \times 5C2 }{13C6 } = 0.4079\] \[\Large \frac{ 5C4 \times 8C2 }{ 13C6 } = 0.08159\]